The lager of two integers is 8 less than twice the smaller. When the smaller number is subtracted from the larger, the difference is 17. Find the two numbers.
Answer by ankor@dixie-net.com(22740)

Two numbers x = the larger, y = the smaller
:
Write an equation for each statement:
:
"The larger of two integers is 8 less than twice the smaller."
x = 2y - 8
:
"When the smaller number is subtracted from the larger, the difference is 17."
x - y = 17
:
Find the two numbers.
;
Using the 1st statement equation, substitute (2y-8) for x in 2nd equation
(2y-8) - y = 17
2y - y = 17 + 8
y = 25
:
Then find x:
x = 2(25) - 8
x = 50 - 8
x = 42
:
Check solution in the statement:
" When the smaller number is subtracted from the larger, the difference is 17."
42 - 25 = 17